find the y-intercept of the line: y+3=4(x+3)
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

find the y-intercept of the line: y+3=4(x+3)
-----
y = 4x+12-3
y = 4x+9
-----
y-intercept: ?
Let x = 0, then y = 9
---
y-intercept (0,9)
==================
Cheers,
Stan H.

Answer by Alan3354(69443)

find the y-intercept of the line: y+3=4(x+3)
------------
The line or function crosses the y=axis at x = 0
sub 0 for x